A First Look at harpoon.sty

Tobias Kuipers

                kuipers@fwi.uva.nl

November 2, 1994

1 Introduction

This packages provides a number of harpoons to be set over or under arbitrary text. The commands are invoked as \someharp{text}, which will put some harpoon over text. text is typeset in LR mode. If you want text to be typeset in math mode you should type \someharp{$text$}.

2 The Commands

The package is invoked with 1 option: \usepackage[xxx]{harpoon}, where xxx is your favourite graphics device driver. On most Unix machines, this will be dvips and on a Macintosh this could be oztex. Refer to the graphics package for more information.

The commands are

• \overleftharp, which looks like ↼ this

• \overrightharp, which looks like ⇀ this

• \overleftharpdown, which looks like ↽ this

• \overrightharpdown, which looks like ⇁ this

• \underleftharp, which looks like ↼ this

• \underrightharp, which looks like ⇀ this

• \underleftharpdown, which looks like ↽ this

• \underrightharpdown, which looks like ⇁ this